OpenStudy (anonymous):

I need help with this algebra question? 13A - 8Z - 2A - 12Z

OpenStudy (anonymous):

13A and -2A are like terms, add -2A to 13A to get 11A. \[11A-8Z-12Z\] Since -8Z and -12Z are like terms, subtract 12Z from -8Z to get -20Z. \[11A-20Z\]

OpenStudy (anonymous):

Thanks Austin

OpenStudy (anonymous):

Need anymore help @Cammie19 ?

OpenStudy (anonymous):

4 5/8 divided by 2 3/4

OpenStudy (anonymous):

Convert the mixed number 4 5/8 into a improper fraction first by multiplying the denominator (8) by the whole number part (4) and add the numerator(6) to get the new numerator. Place new numerator (37) over the old denominator(8). \[\frac{ 37 }{ 8\div2, \frac{ 3 }{ 4 } }\] Convert the mixed number 2 3/4 into an improper fraction first by multiplying the denominator (4) by the whole number part (2) and add the number (3) to get the new numerator. Place the new numerator (11) over the old denominator (4). \[\frac{ 37 }{ 8\div (\frac{ 11 }{ 4 }) }\] \[(\frac{ 37 }{ 8 })(\frac{ 4 }{ 11 })\] Cancel the common factor of 4. \[(\frac{ 37 }{ 2})(\frac{ 1 }{ 11 })\] Simplify \[\frac{ 37 }{ 22 }\] Exact Form: \[\frac{ 37 }{ 22 }\] Decimal Form: 1.68 Mixed Number Form: \[1\frac{ 15 }{ 22}\]

OpenStudy (anonymous):

@Cammie19

OpenStudy (anonymous):

Need anymore help?

OpenStudy (anonymous):

no thanks I'm done and thans again

OpenStudy (anonymous):

4 2/3 divided by 1 1/4

OpenStudy (anonymous):

is anybody going to help me? Austin1617 help me please